OpenStudy (anonymous):

Luan stores coffee beans in a steel container with a capacity of 323.4 cubic inches. He uses a glass container to store sugar. The two containers are congruent cylinders. If the glass container has a height of 8.4 inches, what is its radius?

OpenStudy (anonymous):

3.5 inches 4.2 inches 1.2 inches 6.2 inches

OpenStudy (anonymous):

@AccessDenied @Directrix

Directrix (directrix):

The volume of a right circular cylinder is V = pi r^2 h. 323.4 = pi * r^2 * 8.4 38.5 = pi * r^2 12.262 = r^2 r = 3.50

OpenStudy (anonymous):

I understand now @Directrix YUR THE BEST